I have been stuck on this problem all day, please explain:
Under normal conditions a population of rabbits has an exponential growth rate of 11.8% per day. Consider an initial population of 100 rabbits.
A. Find the exponential growth function. B. When will the population reach 1,000 rabbits? C. How many rabbits will there be after 30 days?

Answer by ewatrrr(24785)

 
Hi,
B(t) = 100(1.118)^t
1000/100 = (1.118)^t t = time in days
log( 10)/log(1.118) = t = 20.64 days
B(30) = 100(1.118)^30 = 2840